500 students are enrolled in at least two of these three classes: Math, English, and History. 170 are enrolled in both Math and English, 150 are enrolled in both History and English, and 300 are enrolled in Math and History. How many of the 500 students are enrolled in all three?

Found 2 solutions by scott8148, ankor@dixie-net.com:
Answer by scott8148(6628)

let x="students enrolled in all three"

"170 are enrolled in both Math and English" __ so 170-x are enrolled in ONLY Math and English

"150 are enrolled in both History and English" __ so 150-x are enrolled in ONLY History and English

"300 are enrolled in Math and History" __ so 300-x are enrolled in ONLY Math and History

"500 students are enrolled in at least two of these three classes"
so (170-x)+(150-x)+(300-x)+x = 500

620-2x=500

120=2x

60=x

Answer by ankor@dixie-net.com(22740)

500 students are enrolled in at least two of these three classes: Math,
English, and History. 170 are enrolled in both Math and English,
150 are enrolled in both History and English, and 300 are enrolled in
Math and History.
How many of the 500 students are enrolled in all three?
:
Let x = no. students enrolled in all three
:
These students are represented in an equal amt in all three groups:
:
The equation for students taking only two subjects
(170-x) + (150-x) + (300-x) = 500 - x
620 - 3x + x = 500
-2x = 500 - 620
-2x = -120
x = +60 students take all three subjects
:
Check solution
170 - 60 = 110 students taking Math & Eng only
150 - 60 = 90 students taking Hist & Eng only
300 - 60 = 240 students taking Math & Hist only
All 3 sub = 60 students
----------------
Tot stud =500
;
: